# NIR-Assisted Image Denoising: A Selective Fusion Approach and A Real-World Benchmark Dataset

### v1

@@ VI-B Comparisons with More Fusion Manners
We compare our selective fusion manner with other fusion methods, including feature summation, channel attention [51], and cross attention [52] to demonstrate our method’s superiority. The results on Real-NAID dataset are shown in Table V. It can be observed the usage of NIR images and other prevailing fusion methods can improve performance to some degree. However, these methods do not achieve competitive performance compared with our SFM while sometimes introducing higher computation costs. It illustrates the superiority of the proposed method.
@@ VI-C Effect of SFM
Effect of LMM. As shown in Table VI, there is a 0.1dB drop on PSNR when removing LMM from SFM. We provide a qualitative result to demonstrate its impact on visual results. As shown in Fig. 7 (b), the incorporation of our LMM in SFM helps fine-scale texture recovery. Moreover, we explore the appropriate receptive field size in LMM by employing varying kernel sizes of DWConv as shown in Table VII. Generally, a larger kernel size leads to greater performance improvement, which shows that a large reception field helps the local modulation of features. But it is improved marginally when the kernel size is larger than $5\times 5$. For the sake of simplicity and efficiency, we set the kernel size of DWConv to $5\times 5$ as default.
Effect of GMM and LMM Arrangement. Here, we conduct experiments with different arrangements of GMM and LMM. The results are shown in Table VIII. ‘GMM + GMM’ and ‘LMM + LMM’ mean we modulate features with 2 GMMs and 2 LMMs respectively. They result in limited performance gain. It shows that our performance improvement is not due to a simple increase in parameter numbers. ‘LMM + GMM’ means we modulate features first locally and then globally. It also leads to limited improvement. It may be because the significant difference in global content leads to inaccurate local feature modulation. Therefore, we deploy a GMM to handle color discrepancy first followed by an LMM dealing with structure discrepancy, dubbed ‘GMM + LMM’, achieving better results.
Effect of Number of SFM. Here we investigate the effect of incorporating multiple SFMs at each scale of NIR-NAFNet. The results are shown in Table IX. It can be observed that the performance generally increases marginally as the number of SFMs grows. Also for the sake of simplicity and efficiency, we only set the number of SFM to 1 at each scale.
@@
TABLE IX: Quantitative comparison of different numbers of SFM.
@@ VI-D Effect of Multi-Scale Loss
The multi-scale loss calculates the difference between output and supervision at different scales, and it can slightly improve performance with few increasing training costs. Taking NAFNet [24] as an example, we evaluate its effect by comparing it with naive $\ell_{2}$ loss. Table X shows that it improves the performance of both single-image and NIR-assisted image denoising. And the improvement is more obvious for the latter. This shows that multi-scale loss can help SFM work better at each scale.
TABLE X: Comparisons of different loss functions.
@@
VII Conclusion
@@ Acknowledgments
This work was supported by the National Natural Science Foundation of China (NSFC) under Grants No. U19A2073.

### v4

@@ VI-B Comparisons with More Fusion Manners
We compare our selective fusion manner with other fusion methods, including feature summation, channel attention [55], and cross attention [56] to demonstrate our method’s superiority. The results on Real-NAID dataset are shown in Table V. It can be observed the usage of NIR images and other prevailing fusion methods can improve performance to some degree. However, these methods do not achieve competitive performance compared with our SFM while sometimes introducing higher computation costs. This may be due to methods like MDTA [10] or cross attention [56] only focus on correlation on the channel dimension or spatial dimension, while our proposed method focuses on both dimensions. These illustrate the superiority of the proposed method.
@@ VI-C Effect of SFM
Figure 12: Visual comparison of different kernel sizes of the DWConvs in our LMM.
@@
Effect of LMM. As shown in Table VI, there is a 0.1dB drop on PSNR when removing LMM from SFM. We provide a qualitative result to demonstrate its impact on visual results. As shown in Fig. 11 (b), the incorporation of our LMM in SFM helps fine-scale texture recovery. Moreover, we explore the appropriate receptive field size in LMM by employing varying kernel sizes of DWConv as shown in Table VII. Larger convolution kernels have larger receptive fields, thus it is beneficial for local feature extraction [57]. Then, better local feature extraction is conducive to better prediction of the fusion weights of NIR and RGB features, thus making better use of NIR image information. Finally, in terms of visual effects, it helps better structure recovery, as shown in Fig. 12. But it is improved marginally when the kernel size is larger than $5\times 5$. For the sake of simplicity and efficiency, we set the kernel size of DWConv to $5\times 5$ as default.
TABLE IX: Quantitative comparison of two GMMs or LMMs in one SFM.
[…]
1 1 25.26 / 0.7695 / 0.299
2 1 25.26 / 0.7691 / 0.301
1 2 25.26 / 0.7715 / 0.300
2 2 25.27 / 0.7711 / 0.299
Effect of GMM and LMM Arrangement. Here, we conduct experiments with different arrangements of GMM and LMM. The results are shown in Table VIII. ‘GMM + GMM’ and ‘LMM + LMM’ mean we modulate features with 2 GMMs and 2 LMMs respectively. They result in limited performance gain. It shows that our performance improvement is not due to a simple increase in parameter numbers. ‘LMM + GMM’ means we modulate features first locally and then globally. It also leads to limited improvement. It may be because the significant difference in global content leads to inaccurate local feature modulation. Therefore, we deploy a GMM to handle color discrepancy first followed by an LMM dealing with structure discrepancy, dubbed ‘GMM + LMM’, achieving better results. Additionally, we conduct experiments on different numbers of GMM and LMM in one SFM. Table IX shows more GMMs or LMMs in one SFM only get minor improvement. For simplicity, We set one GMM and one LMM in one SFM.
Effect of Numbers of Blocks in GMM and LMM. We conduct experiments on the effect of the number of blocks in Table X. Table X shows that more blocks in GMM and LMM can both improve performance. However, increasing the number of $1\times 1$ and $5\times 5$ blocks from one to five only brings a 0.04 dB PSNR improvement, which is minor. For the sake of simplicity and efficiency, we only adopt one $1\times 1$ block in GMM and one $5\times 5$ block in LMM.
Effect of Number of SFM. Here we investigate the effect of incorporating multiple SFMs at each scale of NIR-NAFNet. The results are shown in Table XI. It can be observed that the performance generally increases marginally as the number of SFMs grows. Also for the sake of simplicity and efficiency, we only set the number of SFM to 1 at each scale.
@@
TABLE X: Quantitative comparison of different numbers of blocks in GMM and LMM.
[…]
1 1 25.26 / 0.7695 / 0.299
3 1 25.27 / 0.7702 / 0.302
5 1 25.27 / 0.7701 / 0.302
1 3 25.27 / 0.7709 / 0.299
1 5 25.28 / 0.7701 / 0.302
3 3 25.28 / 0.7705 / 0.300
5 5 25.30 / 0.7718 / 0.302
TABLE XI: Quantitative comparison of different numbers of SFM.
@@ VI-D Effect of Multi-Scale Loss
The multi-scale loss calculates the difference between output and supervision at different scales, and it can slightly improve performance with few increasing training costs. Note that we conduct the experiment several times to eliminate randomness. The results are consistent, only with a small variance. Taking NAFNet [27] as an example, we evaluate its effect by comparing it with naive $\ell_{2}$ loss. Table XII shows that it improves the performance of both single-image and NIR-assisted image denoising. And the improvement is more obvious for the latter. This shows that multi-scale loss can help SFM work better at each scale.
TABLE XII: Comparisons of different loss functions.
@@
VII Limitations
Denoising models in this work are only trained on images captured by one camera. Given that the noise models of different cameras may be different, our denoising models may not generalize well on other cameras or devices. In the future, we consider constructing a larger real-world NIR-RGB dataset covering multiple cameras. Thus, denoising models trained on this dataset may generalize well.
VIII Conclusion
@@ Acknowledgments
This work was supported by the National Natural Science Foundation of China (NSFC) under Grants No. U22B2035.